### v6

# Morphisms from a very general hypersurface

Abstract.
Let $X$ be a very general hypersurface of degree $d$ in the projective $(n+1)$-space with $n\geq 3$, and $f:X\to Y$ a non-birational surjective morphism to a normal projective variety $Y$. We first prove that $Y$ is a klt Fano variety if $\deg f\geq C$ for some constant $C=C(n,d)$ depending only on $n$ and $d$. Next we prove an optimal upper bound $\deg f\leq\deg X$ provided that $Y$ is factorial, $\deg f$ is prime and $\deg f\geq E(n)$ for some constant $E(n)$ (with $E(n)=n(n+1)$ when $Y$ is smooth). As a corollary, we show that $Y\cong{\mathbb{P}}^{n}$ under some conditions on $Y$ and $\deg f$.
Key words and phrases:
Dominant rational map, Hypersurface, Fano variety, Rationally connected
2010 Mathematics Subject Classification:
Primary 14E05;  Secondary 14D05,14J45,14J70,14M22
Dedicated to James McKernan on the occasion of his 60th birthday.
1. Introduction
We work over the field $\mathbb{C}$ of complex numbers.
To motivate our result, we recall that if $X$ is a smooth projective variety of general type, then the set of dominant rational maps $X\dashrightarrow Y$ from $X$ to a smooth variety $Y$ of general type, up to birational equivalence of $Y$, is a finite set. The proof was given by Maehara [32] under the assumption of effective birationality of pluri-canonical maps of varieties of general type. This effective birationality was proved later (cf. Hacon and McKernan [18, Theorem 1.1, Corollary 1.4], Takayama [38], and Tsuji [39]).
If we allow the co-domain $Y$ to be uniruled (or even rational), then the above finiteness result is not true since every variety has a generically finite map to a projective space while the latter has self maps of unbounded degrees. In this paper, we assume that the domain $X$ is a very general hypersurface.
In connection with the famous conjecture:
Let $Y$ be a Fano manifold of Picard number 1. Suppose that $Y$ admits a non-isomorphic surjective endomorphism. Then $Y$ is a projective space.
which has been proved in dimension $3$ by Amerik [2] and Amerik-Rovinsky-Van de Ven [4], independently by Hwang-Mok [22], and even in arbitrary characteristics, coprime with $\deg f$, by Kawakami-Totaro [23], it is natural to consider the following about morphisms from a very general hypersurface (cf. [2, Conjecture B]):
Conjecture 1.1.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface with $n\geq 2$ and $\operatorname{Pic}(X)={\mathbb{Z}}$ (this holds when $n\geq 3$ or $\deg X\geq 4$), and $f:X\to Y$ a non-birational surjective morphism to a smooth projective variety $Y$ of positive dimension. Then there is a constant $C(X)$ depending on discrete invariants of $X$, such that if $\mathrm{deg}~{}f\geq C(X)$, then $Y$ is isomorphic to ${\mathbb{P}}^{n}$.
Conjecture 1.1 in dimension $\leq 3$ has been successfully proved by Amerik, and more recently, by Kawakami and Totaro. Indeed, see Amerik [3, Theorems 1.1 and 1.2] when $n=2$, or when $n=3$ and $Y$ is neither the Mukai–Umemura variety $V_{22}^{s}$ nor of type $V_{5}$; see [2, Proposition 2.2] for the last two Fano $3$-fold cases. See also [23, Proof of Proposition 2.7, Proposition 3.8] (which works in arbitrary characteristics) and [3, Corollary 2.2, Remark 2.4, Lemma 3.0]. We refer to Shao-Zhong [37] for their results for morphisms between Fano manifolds, and more references and progress towards the famous conjecture and Conjecture 1.1.
Here, we achieve a partial result for Conjecture 1.1. Our first main result (Theorem 1.2) is to show that $Y$ is a klt Fano variety if $\deg f$ is large enough.
Theorem 1.2.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface with $n\geq 3$, and $f:X\to Y$ a non-birational surjective morphism to a normal projective variety $Y$. Then we have:
  1. (1)
    $Y$ has only klt singularities; either the canonical divisor $K_{Y}$ is ample, or $-K_{Y}$ is ample.
  2. (2)
    There exists a constant $C=C(n,d)$ such that if $\mathrm{deg}~{}f\geq C$, then $Y$ is a klt Fano variety.
As remarked earlier, Conjecture 1.1 in dimension $3$ has been confirmed by Amerik and more recently by Kawakami and Totaro, both thanks to the known classification list of smooth Fano $3$-folds due to Iskovskikh and Mori-Mukai. However, a classification of Fano manifolds in higher dimensions is unlikely to be feasible. Therefore, for the morphism $f:X\to Y$ in the above theorem, assuming further that $\deg f$ is large with additional conditions, we prove in Theorem 1.3 and Corollaries 4.7 and 4.8 that $Y\cong{\mathbb{P}}^{n}$, thus confirming Conjecture 1.1. This also gives the precise constant $C(n,d)$ and even removes or weakens the smoothness assumption on $Y$.
Before stating our next main results, we introduce some notation. Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$ with $n\geq 2$, and $f:X\to Y$ a non-isomorphic surjective morphism of degree $m$ to a normal projective variety $Y$. Consider the restriction map $f_{H_{X}}:=f|_{H_{X}}$ where $H_{X}:=H\cap X$ with $H\subset{\mathbb{P}}^{n+1}$ a hyperplane, it follows that $f_{H_{X}}$ is finite. Let $H_{Y}:=f(H_{X})$. Then, by the projection formula, $f_{*}f^{*}H_{Y}=(\deg f)H_{Y}$. Therefore, $\deg f_{H_{X}}$ divides $\deg f$ when $\mathrm{Pic}(X)=\mathbb{Z}$; indeed, writing $f^{*}H_{Y}\sim rH_{X}$ for some positive integer $r$ and using the projection formula to push forward by $f_{*}$, we then have $\deg f=r\deg(f_{H_{X}}:H_{X}\to H_{Y})$. We define a linear subspace $S(f)$ of $|H_{X}|$ below (which is the pullback of a complete linear system on $Y$; see Proposition 4.2)
$S(f):=\{H^{\prime}\in|H_{X}|\,;\,\deg f|_{H^{\prime}}=\deg f\}.$
The $Y$ in Theorem 1.3 below, which is the image of smooth $X$, has at worst klt singularities (cf. Proposition 2.2). So, it is quite natural to further assume that $Y$ has at worst $\varepsilon$-lc singularities (cf. Birkar [8, 2.7] for its definition). Recall that a canonical singularity is $\varepsilon$-lc with $\varepsilon=1$. The $Y$ in Theorem 1.3 or 1.4 below, being klt, has at worst canonical singularities if $Y$ is further factorial. Let $Y_{\rm reg}:=Y\setminus\operatorname{Sing}Y$ be the smooth locus of $Y$.
Theorem 1.3.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$ and $f:X\to Y$ a surjective morphism to a normal projective variety $Y$ with at worst $\varepsilon$-lc singularities. Assume either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$. Assume further the fundamental group $\pi_{1}(Y_{\rm reg})=(1)$ (which holds if $Y$ is smooth or $\deg f$ is prime by Proposition 2.2). Then we have:
  1. (1)
    $\deg f\leq D_{1}d$ for some constant $D_{1}=D_{1}(n,\varepsilon)$, if $-K_{Y}$ is ample and $S(f)\neq\emptyset$.
  2. (2)
    $\deg f\leq{\rm max}\{D_{1}d,D_{2}\}$ for some constant $D_{2}=D_{2}(n,\varepsilon,d)$, if $\deg f$ is prime.
  3. (3)
    If $\deg f\geq d$, then $\deg f=d$ and $Y={\mathbb{P}}^{n}$, unless ${\rm Codim}~{}S(f)\geq 2$ in $|H_{X}|$.
Theorem 1.4 below says that in Theorem 1.3, if we assume $Y$ is factorial then $D_{1}=1$, and $D_{2}$ depends only on $n$. The upper bound in Theorem 1.4 below is optimal because of the projection map (see Remark 1.5).
An upper bound of $\deg f$ without a condition like $\deg f$ being prime or so in Theorem 1.3 or 1.4, seems unlikely. Indeed, a projection $f_{0}:X\to{\mathbb{P}}^{n}$ composed with quotient maps $f_{s}:{\mathbb{P}}^{n}\to Y_{s}:={\mathbb{P}}^{n}/G_{s}$ for a finite subgroup $G_{s}\subset\operatorname{PGL}_{n+1}({\mathbb{C}})$, will give morphisms from $X$ with unbounded degrees, with $Y_{s}$ being smooth (and hence is isomorphic to ${\mathbb{P}}^{n}$) when $G_{s}$ is generated by pseudo-reflections, or being singular otherwise.
Theorem 1.4.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$ and $f:X\to Y$ a surjective morphism to a factorial normal projective variety $Y$. Assume $n\geq 3$, $d\geq 3$, and $\pi_{1}(Y_{\rm reg})=(1)$ (which holds if $Y$ is smooth or $\deg f$ is prime by Proposition 2.2). Then we have:
  1. (1)
    $\deg f\leq d$, if $-K_{Y}$ is ample and $S(f)\neq\emptyset$.
  2. (2)
    $\deg f\,\leq\,{\rm max}\{d,\,\,D_{2}/d^{1/(n-1)}\}$ for some constant $D_{2}=D_{2}(n)$, if $\deg f$ is prime.
  3. (3)
    $\deg f\leq d$, if $\deg f$ is prime and $d\geq D_{2}^{(n-1)/n}$.
As a consequence of Theorems 1.3 and 1.4 (or Theorem 4.5), we show that $Y\cong{\mathbb{P}}^{n}$ under some conditions on $Y$ and $\deg f$ (see Corollaries 4.7 and 4.8).
Remark 1.5.
  1. (1)
    If $f:X\to Y={\mathbb{P}}^{n}$ is a projection morphism from a point in ${\mathbb{P}}^{n+1}\setminus X$, then $\deg f=d=\deg X$ and ${\rm Codim}~{}S(f)=1$ in $|H_{X}|$. Indeed, the pullbacks to $X$ of hyperplanes in $Y$ give $S(f)$. In particular, the upper bound in Theorem 1.4 is optimal. On the other hand, we have $\deg f\geq d-n$, by Lemma 5.4, when $X$ is of general type (i.e., $d\geq n+3$).
  2. (2)
    In Theorem 1.4, we can take $D_{2}=(D_{2}^{\prime})^{1/(n-1)}$ in terms of upper bound $D_{2}^{\prime}$ of $(-K_{Y})^{n}$ (see Theorem 4.5). So the condition in Theorem 1.4 (3) is $d\geq D_{2}^{(n-1)/n}=(D_{2}^{\prime})^{1/n}$ ($=n(n+1)$ when $Y$ is smooth, see [27, Corollary 1]).
  3. (3)
    In Theorem 1.3, when $\deg f=d$, we expect that $f$ is a projection morphism, and hence the exceptional case in Theorem 1.3 (3) is expected not to occur.
  4. (4)
    In Theorem 1.3, when $\deg f\geq d$, ``${\rm Codim}\,S(f)\leq 1$ in $|H_{X}|$” holds if and only if $\deg f=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(f(H_{X}))={\mathcal{O}}(1)$ (cf. Proposition 4.2).
  5. (5)
    In Theorem 1.4, the condition $n\geq 3$ and $d\geq 3$ can be weakened as either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$ (just the same as in Theorem 1.3).
  6. (6)
    From the Lefschetz theory and the irreducibility of the monodromy action on vanishing cycles (cf. Voisin [40, Chapters 1-3]), the non-birational curve image of a very general plane curve of degree $d\neq 3$ is a rational curve. Hence we consider only the case $n\geq 2$ in our theorems.
Our proofs of the theorems combines Hodge theory and some birational geometry including Hacon-McKernan-Xu’s work [19] and Birkar’s work [8]. Fujita’s theory ([13], [14]) on (quasi-) polarized projective varieties and $\Delta$-genus is also crucial in proving Proposition 4.1 (essentially from Höring-Peternell [21, Theorem 3.2]) and Proposition 4.2, which in turn, are needed for proving Theorems 4.5, 1.3 and 1.4; see also Kobayashi-Ochiai [24].
The method presented here can be used to obtain similar results for a very general hypersurface of large degree in a Fano manifold of Picard number one; see Remarks 2.6 and 3.3.
It is not clear at the moment how to treat the case where $f$ is not a morphism.
Acknowledgements. The authors would like to thank IBS and KIAS in Korea, NUS in Singapore, and Simons Center for Geometry and Physics in the United States for their warm hospitality, and E. Amerik, G. Zhong and the referee for the valuable suggestions. Y. Lee is supported by the Institute for Basic Science IBS-R032-D1, Y. Luo by a PTA fellowship of NUS, and D. -Q. Zhang by ARF: A-8002487-00-00 of NUS.
2. Preliminary results
2.1.
Terminology and Notation We adopt the standard terminology and notation as in [20], and push-forward of Cartier divisors for a proper morphism (which preserves linear equivalence) as in [15, Proposition 1.4]. We also refer to [26] or [8, 2.7] for the definitions of Kawamata log terminal (klt), canonical and $\varepsilon$-lc (with $\varepsilon>0$) singularities.
Let $f:X\dashrightarrow Y$ be a dominant rational map between two varieties of the same dimension. The degree of $f$, denoted as $\mathrm{deg}~{}f$, is defined as the degree of the function field extension $K(X)/K(Y)$. It is also equal to the cardinality of $f^{-1}(y)$ for a general point $y\in Y$.
For a finite surjective morphism $f:X\to Y$ between normal projective varieties, we use very often the ramification divisor formula
$K_{X}=f^{*}K_{Y}+R_{f}$
and pullback $f^{*}D$ for a Weil divisor $D$ on $Y$. Precisely, we restrict $D$ on the smooth locus of $Y$, use $f^{*}$ to pull the restriction back, and then take the Zariski closure. This pullback definition coincides with the usual definition when $D$ is a $\mathbb{Q}$-Cartier divisor.
A hypersurface $X\subset{\mathbb{P}}^{n+1}$ of degree $d$ is very general if it is outside the union of countably many proper closed subsets of hypersurfaces of degree $d$.
We begin with the following key proposition.
Proposition 2.2.
(Key Proposition) Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$, $Y$ a normal projective variety of positive dimension, and $f:X\to Y$ a surjective morphism. Suppose either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$. Then the following hold (indeed, (1) - (3) hold whenever either $n\geq 3$ and $d\geq 1$, or $n=2$ and $d\geq 4$).
  1. (1)
    $\operatorname{Aut}(X)=(1)$.
  2. (2)
    Let $H_{X}$ be a general section of ${\mathcal{O}}_{{\mathbb{P}}^{n+1}}(1)$ restricted to $X$. Then $\operatorname{Pic}X={\mathbb{Z}}[H_{X}]$ (so the Picard number $\rho(Y)=1$ and every member of $|H_{X}|$ is reduced and irreducible), and the fundamental group $\pi_{1}(X)=(1)$ (so there is no non-trivial torsion line bundle on $X$).
  3. (3)
    $f$ is a finite morphism,
  4. (4)
    $Y$ is ${\mathbb{Q}}$-factorial and klt.
Suppose further that $f$ is non-birational. Then:
  1. (5)
    If $d\leq n+2$ then $Y$ is a klt Fano variety.
  2. (6)
    $H^{i}(Z,{\mathcal{O}}_{Z})=0$ for any smooth model $Z$ of $Y$ and all $i>0$.
  3. (7)
    $p_{g}(Z)=0$, and $\chi({\mathcal{O}}_{Z})=1$.
  4. (8)
    $\pi_{1}(Z)=(1)=\pi_{1}(Y)$.
  5. (9)
    For $Y_{\mathrm{reg}}:=Y\setminus\operatorname{Sing}Y$, we have $|\pi_{1}(Y_{\mathrm{reg}})|$ being a finite number dividing $\deg f$.
  6. (10)
    Suppose $\deg f$ is prime. Then $\pi_{1}(Y_{\mathrm{reg}})=(1)$. So for Weil integral divisors $M_{j}$ on $Y$ with $sM_{1}\sim sM_{2}$ for some integer $s\geq 1$, we have $M_{1}\sim M_{2}$.
Proof.
By the assumption on $n$ and $d$, we have $\operatorname{Aut}(X)=(1)$ by [33, page 347], and also
$\pi_{1}(X)=\pi_{1}({\mathbb{P}}^{n+1})=(1)$
and the restriction map
$\operatorname{Pic}({\mathbb{P}}^{n+1})\to\operatorname{Pic}(X)$
is an isomorphism by the Lefschetz hyperplane theorem (cf. [34] and [30]). Then (2) follows.
As in (2), write
${\mathbb{Z}}[H_{X}]=\operatorname{Pic}(X)=\operatorname{Pic}(X)/\operatorname{Pic}^{0}(X)=\mathrm{NS}(X)$
where $H_{X}$ is a section of ${\mathcal{O}}_{{\mathbb{P}}^{n+1}}(1)$ restricted to $X$ and it is very ample. Since $\dim\,Y>0$, $\rho(Y)=1=\rho(X)$, hence $f^{*}$ induces an isomorphism
$\mathrm{NS}(Y)_{{\mathbb{Q}}}\to\mathrm{NS}(X)_{{\mathbb{Q}}}$
of ${\mathbb{Q}}$-coefficient Neron-Severi spaces, so $H_{X}$ is equal to some $f^{*}D_{Y}$. Thus (3) follows.
Since $f:X\to Y$ is a finite surjective morphism and $X$ is smooth (and hence ${\mathbb{Q}}$-factorial and klt), we obtain (4) by [26, Lemma 5.16 and Proposition 5.20].
For (9), since $\dim f^{-1}(\operatorname{Sing}Y)$ is of codimension $\geq 2$ in the smooth variety $X$ and by [10, Proposition 1.3], the image of $(1)=\pi_{1}(X)=\pi_{1}(f^{-1}\,Y_{\mathrm{reg}})$ in $\pi_{1}(Y_{\mathrm{reg}})$ is of index $\leq\deg f$, so $|\pi_{1}(Y_{\mathrm{reg}})|$ is finite, and further, $X\to Y$ factors as $X\to Y^{\prime}\to Y$, where $Y^{\prime}\to Y$ is finite and quasi-étale and is the normalisation of $Y$ in the function field of the universal cover of $Y_{\mathrm{reg}}$. Thus $|\pi_{1}(Y_{\mathrm{reg}})|=\deg(Y^{\prime}/Y)$ divides $\deg f$, finishing the proof of (9).
For (10), as in (9), we have $X\to Y^{\prime}\to Y$ with $|\pi_{1}(Y_{\mathrm{reg}})|=\deg(Y^{\prime}/Y)$ dividing $\deg f$. If $|\pi_{1}(Y_{\mathrm{reg}})|\geq 2$ then the primality of $\deg f$ implies that $X=Y^{\prime}$, and hence $\operatorname{Gal}(Y^{\prime}/Y)\leq\operatorname{Aut}(X)=(1)$, which is absurd. Thus $\pi_{1}(Y_{\mathrm{reg}})=(1)$. If $sL\sim 0$ for some integer $s\geq 2$ while $L\not\sim 0$, taking a prime factor $q$ of $s$ and replacing $L$ by a multiple, we may assume that $qL\sim 0$ while $L\not\sim 0$. This relation induces a finite Galois ${\mathbb{Z}}/q{\mathbb{Z}}$-cover $Y^{\prime}\to Y$, which is quasi-étale, i.e., étale over $Y_{\mathrm{reg}}$ by the purity of branch loci. Thus $|\pi_{1}(Y_{\mathrm{reg}})|\geq q$, contradicting its triviality which we have just proved. This proves (10).
Next we prove (5) - (8). By the ramification divisor formula,
$K_{X}=f^{*}K_{Y}+R_{f}$
with $R_{f}$ the ramification divisor. Since $\rho(X)=1$ and $\pi_{1}(X)=(1)$, we have linear equivalence
$f^{*}K_{Y}\sim k_{y}H_{X}$
for some integer $k_{y}$.
Claim 2.3.
It is impossible that $k_{Y}=0$ and $R_{f}=0$.
We prove the claim. Suppose the contrary that $k_{y}=0$ and $R_{f}=0$. Then $K_{X}\sim 0$, and $f$ is quasi-étale. Let
$X^{\prime}\to X\to Y$
be the Galois closure of $f$. Then $X^{\prime}\to X$, like $f$, is still quasi-étale, and hence étale since $X$ is smooth and by the purity of branch loci. Thus $X^{\prime}=X$ because $\pi_{1}(X)=(1)$. So $f:X\to Y$ is Galois with Galois group $G$. Now $G\leq\operatorname{Aut}(X)=(1)$. Hence $\deg f=|G|=1$, contradicting the non-birationality assumption of $f$. This proves Claim 2.3.
We return back to the proof of the proposition.
If $k_{y}<0$, then $-K_{Y}$ is ample. Thus $Y$ is a klt Fano variety, and hence both $Y$ and its smooth model $Z$ are rationally connected, so (6) - (7) clearly hold in this case. Since smooth and hence normal rationally connected varieties are simply connected [10, Theorem 3.5], the assertion (8) also holds in this case.
Thus, it remains to show (5) - (8) under the assumption that either $k_{y}>0$, or $k_{y}=0$ and $R_{f}\neq 0$. In either case, from the ramification divisor formula we see that
$K_{X}\sim(d-n-2)H_{X}$
is a nonzero effective divisor, and hence ample since $\rho(X)=1$. Thus $d\geq n+3\geq 5$ (this also proves (5)). Further, $K_{X}$ is very ample and $X$ is of general type. Replacing $Y$ by its smooth model $Z$ and $X$ with a higher smooth birational model $W$, our $f$ induces a generically finite surjective morphism
$h:W\to Z.$
Since $X$ is simply connected, so is $W$, hence $|\pi_{1}(Z)|=|\pi_{1}(Z):h(\pi_{1}(W))|$ is bounded by the number of connected components of a general fibre, i.e., by $\deg f$ (cf. [10, Proposition 1.3]). Let
$Z^{\prime}\to Z$
be the universal cover. Then $W\to Z$ factors as
$W\to Z^{\prime}\to Z$
because $W$ is simply connected.
Since $d\geq n+3$, by [3, Proposition 2.1 ] or [6, Remark 3.3, Example 1.2], either $p_{g}(Z)=0$, or $p_{g}(Z)=p_{g}(W)$ ($=p_{g}(X)$).
Claim 2.4.
It is impossible that $p_{g}(Z)=p_{g}(W)$ ($=p_{g}(X)$).
Suppose to the contrary that $p_{g}(Z)=p_{g}(W)$. Now the natural injection, via $h^{*}$,
$H^{0}(Z,K_{Z})\to H^{0}(W,K_{W})=H^{0}(X,K_{X})$
is an isomorphism, so the canonical rational map $\Phi_{|K_{W}|}$ (which is birational, since $K_{X}$ is very ample) factors as the composition of rational maps
$W\to Z\overset{\Phi_{|K_{Z}|}}{\dashrightarrow}\Phi_{|K_{Z}|}(Z),$
or equivalently as the composition of rational maps
[…]
where $W\to X$ is birational and $Y\dasharrow Z$ is the inverse of the birational map $Z\dasharrow Y$. Hence $\deg f=1$, a contradiction. This proves Claim 2.4.
We return back to the proof of the proposition. By Claim 2.4, $\dim H^{n}(Z,{\mathcal{O}}_{Z})=p_{g}(Z)=0$. Applying the Kodaira vanishing to the cohomologies of the exact sequence
$0\to{\mathcal{O}}_{{\mathbb{P}}^{n+1}}(-X)\to{\mathcal{O}}_{{\mathbb{P}}^{n+1}}\to{\mathcal{O}}_{X}\to 0$
we have $H^{i}(X,{\mathcal{O}}_{X})=0$ for all $i=1,\dots,n-1$, so the same holds for $Z^{\prime}$ and $Z$, since the latter ones are dominated by $X$. Combining the above all, we get $\chi({\mathcal{O}}_{Z})=1$. This proves (6) - (7) in this case too.
Claim 2.5.
$p_{g}(Z^{\prime})=0$.
We prove the claim. Suppose the contrary that $p_{g}(Z^{\prime})\geq 1$. Then $\deg(Z^{\prime}/Z)\geq 2$. Applying [3, Proposition 2.1 ] or [6, Remark 3.3, Example 1.2] and arguing as above, we have
$p_{g}(Z^{\prime})=p_{g}(W)=p_{g}(X)$
and $W\to Z^{\prime}$ is a birational morphism. Now $G=\operatorname{Gal}(Z^{\prime}/Z)$ acts birationally on $X$ and hence biregularly on the canonical model of $X$, which is $X$ itself. Thus $G\leq\operatorname{Aut}(X)=(1)$. Hence $Z^{\prime}=Z$, a contradiction. This proves Claim 2.5
We return back to the proof of part (8) of the proposition. By Claim 2.5, $p_{g}(Z^{\prime})=0$. So $\chi({\mathcal{O}}_{Z^{\prime}})=1$. Since $Z^{\prime}\to Z$ is étale, $\deg(Z^{\prime}/Z)|\,\chi(\mathcal{O}_{Z})|=|\chi({\mathcal{O}}_{Z^{\prime}})|$ ($=1$). Hence $\deg(Z^{\prime}/Z)=1$, i.e., $Z^{\prime}=Z$. Thus $Z$ is simply connected. The same holds for $Y$ since we may assume $Z\to Y$ is a surjective birational morphism and $Y$ is normal (cf. [10, Proposition 1.3]). This proves (8) in this case too. ∎
Remark 2.6.
All assertions in Proposition 2.2, except possibly (4), (5), (8) and (10), can be generalized from ${\mathbb{P}}^{n+1}$ to any nondegenerate Fano manifold $P$ of $\dim\,P\geq 3$ with the Picard number $\rho(P)=1$, if $X$ is a very general hypersurface section of $P$ and $X$ is of general type. If we further assume that $\operatorname{Aut}(X)=(1)$, then all the assertions in Proposition 2.2 can be generalized except (5).
Indeed, let $f:X\to Y$ be a non-birational surjective morphism to a normal projective variety $Y$ of positive dimension. Assuming that $X$ is of general type and $P$ is a Fano manifold with $\rho(P)=1$, we can apply again the irreducibility of the monodromy action on vanishing cycles (cf. [40, Chapters 1-3]) and Kodaira vanishing to show $H^{i}({\mathcal{O}}_{Z})=0$ for any smooth model $Z$ of $Y$ and all $i>0$. Other proofs are basically the same.
3. Proof of Theorem 1.2
Proof of Theorem 1.2.
We may assume that $\mathrm{dim}~{}Y>0$. By Proposition 2.2, $f$ is finite (so $m:=\deg\,f\geq 2$), and $Y$ is ${\mathbb{Q}}$-factorial and klt. We use the notation in Proposition 2.2:
${\rm Pic}\,X={\mathbb{Z}}[H_{X}],\ \ \ f^{*}K_{Y}\sim k_{y}H_{X}.$
If $k_{y}<0$, then $Y$ is a klt Fano variety, and the theorem holds. Thus, we may assume $k_{y}\geq 0$.
Claim 3.1.
$K_{Y}$ is a ${\mathbb{Q}}$-Cartier ample divisor.
The above claim would complete the first part of the theorem. Now we prove this claim.
If $k_{y}>0$, then $f^{*}K_{Y}$ and hence $K_{Y}$ are ample $\mathbb{Q}$-Cartier divisors. Thus we may assume $k_{y}=0$. By the projection formula,
$mK_{Y}=f_{*}f^{*}K_{Y}\sim f_{*}(k_{y}H_{X})=0.$
Let $Y^{\prime}\to Y$ be the index-1 cover which is quasi-étale, so that $K_{Y^{\prime}}\sim 0$ and $Y^{\prime}$ has only canonical singularities. As in the proof of Proposition 2.2, $\pi_{1}(X)=(1)$ implies that $X\to Y$ factors as
$X\to Y^{\prime}\to Y.$
Let $Y^{\prime\prime}$ be a smooth model of $Y^{\prime}$. Then the Kodaira dimension $\kappa(X)\geq\kappa(Y^{\prime\prime})=0$. Hence $K_{X}\sim(d-n-2)H_{X}$ is effective, so $d\geq n+2\geq 5$.
Since $p_{g}(Y^{\prime\prime})=1>0$, the morphism $X\to Y^{\prime}$ is birational, or else contradicting Proposition 2.2 applied to $X\to Y^{\prime}$. Hence $\kappa(X)=\kappa(Y^{\prime\prime})=0$. Thus $d=n+2$ and $K_{X}\sim 0$. So $f:X\to Y$ is quasi-étale. As in Proposition 2.2, $\pi_{1}(X)=(1)$ implies that the Galois closure of $f$ is $f$ itself. Then $\operatorname{Gal}(X/Y)\leq\operatorname{Aut}(X)=(1)$ by Proposition 2.2, which implies that $\deg f=1$, a contradiction. This finishes the proof of Claim 3.1.
We return back to the proof of the theorem. Now $K_{Y}$ is ample. Since $\rho(X)=1$,
$(d-n-2)H_{X}\sim K_{X}=f^{*}K_{Y}+R_{f}$
is ample too. So $d\geq n+3$ and $X$ is of general type.
By Hacon-McKernan-Xu [19, Theorem 1.3] (or Birkar-Zhang [9, Theorem 1.3]), there is a constant $C_{1}=C_{1}(n)$ such that $C_{1}K_{Y}$ is birational, so $C_{1}K_{Y}>J_{Y}$ for some nef and big pullback $J_{Y}$ of an ample Cartier divisor via some pluri-canonical morphism of $Y$. Thus $C_{1}K_{X}>f^{*}J_{Y}$. Since $K_{X}\sim(d-n-2)H_{X}$ and $H_{X}^{n}=d$, if we set
$C:=C(n,d)=(C_{1}K_{X})^{n}$
we have
$C=dC_{1}^{n}(d-n-2)^{n}.$
Now $C_{1}K_{X}>f^{*}J_{Y}$ and the claim below imply
$C=(C_{1}K_{X})^{n}\geq(\deg\,f)J_{Y}^{n}\geq\deg\,f.$
Claim 3.2.
Let $V$ be a projective variety of dimension $n\geq 1$, with $\mathbb{Q}$-Cartier nef divisors $L_{1},L_{2}$. Suppose $M:=L_{1}-L_{2}$ is pseudo-effective. Then
$L_{1}^{n}\geq L_{2}^{n}.$
We prove the claim. The nefness of $L_{j}$ implies
$L_{1}^{n}=L_{1}^{n-1}(L_{2}+M)\geq L_{1}^{n-1}L_{2}=L_{1}^{n-2}(L_{2}+M)L_{2}\geq L_{1}^{n-2}L_{2}^{2}\geq\cdots\geq L_{2}^{n}.$
This proves the claim and also the theorem. ∎
Remark 3.3.
Theorem 1.2 can be generalized from ${\mathbb{P}}^{n+1}$ to any nondegenerate Fano manifold $P$ of $\dim\,P\geq 4$ with the Picard number $\rho(P)=1$, if $X$ is a very general hypersurface section of $P$ and $X$ is of general type. Indeed, let $f:X\to Y$ be a non-birational surjective morphism to a normal projective variety $Y$ of positive dimension. Then there is a constant $C=C(n,d)$ where $d$ is the degree of the hypersurface section $X$ such that if $\deg\,f\geq C$ then $Y$ is a klt Fano variety. The proof is the same as above by Remark 2.6 and by Hacon-McKernan-Xu [19, Theorem 1.3] (or Birkar-Zhang [9, Theorem 1.3]).
Remark 3.4.
Assume $Y$ is smooth in Theorem 1.2. If $d\geq 3$, then Amerik’s result [3, Corollary 2.2] implies that the Hodge numbers of $Y$ coincide either with that of $X$ or with that of ${\mathbb{P}}^{n}$. This excludes some rational homogeneous spaces $G/P$ as $Y$ with $\rho(G/P)=1$ [7, Proposition 5.1]. Again, if $d\geq 3$, then $Y$ can not be a hypersurface of degree $\geq 3$ in ${\mathbb{P}}^{n+1}$, by [11, Corollary 1.1]. Also if $d\geq n+3$ then there exists a constant $C(X)$ depending on discrete invariants of $X$ such that $Y$ can not be a smooth $n$-quadric hypersurface when $\deg f\geq C(X)$ [3, Page 264].
4. Bounds of degrees of morphisms and conditions to be the projective space
The following is essentially from [21] and needed for the proof of Proposition 4.2. Here we have weakened the seemingly needed ampleness of the divisor $D$ omitted there, and used Khovanskii-Teissier equality case and Fujita’s $\Delta$-genus theory for singular varieties to remove the klt assumption on $X$ there and slightly simplified the argument.
For a Weil divisor $D$ on a normal variety $X$, we use ${\mathcal{O}}_{X}(D)$ to denote the corresponding reflexive sheaf; see [36, Proposition (2) and Theorem (3), page 282] for its properties. By abuse of notation, sometimes we use $H^{0}(X,D)$ to denote $H^{0}(X,{\mathcal{O}}_{X}(D))$.
Proposition 4.1.
(cf. [21, Theorem 3.2, Corollary 3.4]) Let $X$ be a normal projective variety of dimension $n$ and $D$ an integral ${\mathbb{Q}}$-Cartier nef and big divisor such that $\dim H^{0}(X,{\mathcal{O}}_{X}(D))\geq n+D^{n}$, and the rational map $\Phi_{|D|}$ is generically finite. Then:
  1. (1)
    Fujita’s $\Delta$-genus $\Delta(X,D):=n+D^{n}-\dim H^{0}(X,D)$ vanishes. There are birational morphisms $\mu:V\to X$ and $h:V\to W$ between normal projective varieties with $V$ being smooth, and very ample divisor $N$ on $W$, such that $\mu^{*}D=h^{*}N$, so $D$ is Cartier and $|D|$ is base point free.
  2. (2)
    If $D$ is ample, then $h$ can be identified with $\mu$ (and $W=X$).
  3. (3)
    If $D$ is ample and $D^{n}\leq 1$, then $X={\mathbb{P}}^{n}$ and ${\mathcal{O}}_{X}(D)={\mathcal{O}}(1)$.
Proof.
By Hironaka’s resolution theorem, there is a birational morphism $\mu:V\to X$ such that $V$ is smooth projective,
$\mu^{*}D=M+E$
where $|M|$ (the moving part) is base point free with $H^{0}(V,M)\cong H^{0}(X,D)$, and $E$ is effective. Since $|D|$ gives a generically finite map by the assumption, the base point free $|M|$ gives a generically finite morphism; thus $M$ is big and nef. By the nefness of $\mu^{*}D$ and $M$ and the effectivity of $E$, we have
(*) $D^{n}=(\mu^{*}D)^{n}\geq(\mu^{*}D)^{n-1}M\geq\cdots\geq M^{n}$
Now the non-negativity of Fujita’s $\Delta$-genus (see [14, Theorem 1.1]) and our assumption imply
$0\leq\Delta(V,M)=n+M^{n}-\dim H^{0}(V,M)=n+M^{n}-\dim H^{0}(X,D)\leq n+D^{n}-\dim H^{0}(X,D)\leq 0$
Thus the above and all the ones in the (*4) are equalities. Hence $\Delta(V,M)=0=\Delta(X,D)$, and
$(\mu^{*}D)^{n-1}.M$ $=M^{n}=D^{n},$
$((\mu^{*}D)^{n-1}M)^{n}$ $=(\mu^{*}D^{n})^{(n-1)}M^{n}.$
Now the last equality case of Kovanskii-Tessier implies that $M$ and $\mu^{*}D$ (and hence also $E$) are proportional in the ${\mathbb{R}}$-coefficient Néron-Severi space (cf. [31, Proposition 3.5]).
If $E\neq 0$ then, being proportional to $M$, is also nef and big, so the first inequality in the (*4) above is strict, which is a contradiction. Thus $E=0$. Hence $\mu^{*}D=M$ is Cartier, so is $D=\mu_{*}M$. On the other hand, by the vanishing of $\Delta(V,M)$, we have
$M=h^{*}N$
for some birational morphism $h:V\to W$ to a normal projective variety and a very ample divisor $N$ (cf. [14, Theorem 1.1]). This proves the assertion (1).
The assertion (2) is clear by the rigidity lemma and ampleness of $D$ and $N$ (cf. [12, Lemma 1.15]). For (3), (1) and (2) and the push-forward by $h=\mu$ imply (on $V=X$)
$D=N,$
so $D$ is integral, Cartier and ample. Now the assumption $D^{n}\leq 1$ means $D^{n}=1$. Hence $(X,{\mathcal{O}}(D))=({\mathbb{P}}^{n},{\mathcal{O}}(1))$ by [13, Corollary 4.3] or [24, Theorem 1.1]. ∎
Below is another key proposition which is needed for Theorems 4.5, 1.3, and 1.4.
Proposition 4.2.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$ and $f:X\to Y$ a surjective morphism of degree $m\geq 2$ to a normal projective variety $Y$. Assume either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$. Assume additionally $\pi_{1}(Y_{\rm reg})=(1)$ (which holds whenever $Y$ is smooth or $m:=\deg f$ is prime by Proposition 2.2). Write $\operatorname{Pic}(X)={\mathbb{Z}}[H_{X}]$ with $H_{X}$ a general section of ${\mathcal{O}}_{{\mathbb{P}}^{n+1}}(1)|_{X}$ as in Proposition 2.2. Let $L_{Y}$ be an integral divisor on $Y$ such that
$f^{*}L_{Y}\sim t_{y}H_{X}$
for some minimal integer $t_{y}\geq 1$ (so that $L_{Y}$ is ample). Then:
  1. (1)
    For every integral Weil divisor $M$ on $Y$ we have $M\sim sL_{Y}$ for some integer $s$.
  2. (2)
    Suppose $Y$ has only $\varepsilon$-lc singularities and $-K_{Y}$ is ample. If $t_{y}=1$, i.e., $f^{*}L_{Y}\sim H_{X}$, then $\deg f\leq D_{1}d$ for some constant $D_{1}=D_{1}(n,\varepsilon)$ (with $D_{1}=1$ when $L_{Y}^{n}$ is an integer).
  3. (3)
    Suppose that $S(f):=\{H\in|H_{X}|\,;\,\deg f|_{H}=\deg f\}$ is non-empty so that $f_{*}H=mH_{Y}$ where $H_{Y}=f(H)$ for some $H\in S(f)$. Then $H_{Y}\sim L_{Y}$, $t_{y}=1$ and $S(f)=f^{*}|L_{Y}|$ (a linear subspace of $|H_{X}|$).
  4. (4)
    $S(f)\neq\emptyset$ holds if and only if $t_{y}=1$ and $|L_{Y}|\not=\emptyset$.
  5. (5)
    Suppose that $m=\deg f\geq d$. Then “ ${\rm Codim}\,S(f)\leq 1$ in $|H_{X}|$” holds if and only if $\deg f=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$.
Proof.
(1) Write $f^{*}M\sim sH_{X}$ with integer $s\neq 0$, say $s\geq 1$ after replacing $M$ by $-M$ if needed. By the minimality of $L_{Y}$, we have $t_{y}|s$. So $f^{*}M\sim f^{*}(s/t_{y})L_{Y}$. Using the projection formula to push forward by $f_{*}$, our $M$ is ${\mathbb{Q}}$-linearly equivalent and hence linearly equivalent to $(s/t_{y})L_{Y}$ by the assumption $\pi_{1}(Y_{\mathrm{reg}})=(1)$ (see Proposition 2.2 (10)).
(2) By Birkar [8, Theorem 1.1], there is a constant $D_{1}^{\prime}=D_{1}^{\prime}(n,\varepsilon)$ such that $|D_{1}^{\prime}L_{Y}|$ gives a birational map, so $(D_{1}^{\prime}L_{Y})^{n}\geq 1$. Taking top intersection, we get $d=H_{X}^{n}=(f^{*}L_{Y})^{n}=(\deg f)L_{Y}^{n}\geq(\deg f)/D_{1}$, where $D_{1}=D_{1}(n,\varepsilon):=(D_{1}^{\prime})^{n}$.
(3) By Proposition 2.2, $\operatorname{Pic}(X)={\mathbb{Z}}[H]={\mathbb{Z}}[H_{X}]$, and every member of $|H_{X}|$ is reduced and irreducible. By the projection formula and the assumption, we have $f^{*}H_{Y}=H\sim H_{X}$. Thus $H_{Y}\sim L_{Y}$ and $t_{y}=1$ (cf. (1)). This shows $S(f)\subseteq f^{*}|L_{Y}|$. Conversely, suppose $H^{\prime}=f^{*}H_{Y}^{\prime}$ ($\sim f^{*}L_{Y}\sim f^{*}H_{Y}=H\sim H_{X}$) for some $H_{Y}^{\prime}\sim L_{Y}$. By the projection formula, $f_{*}H^{\prime}=mH_{Y}^{\prime}$, and hence $H_{Y}^{\prime}$ is reduced (and irreducible) since $m=\deg f$. So $H_{Y}^{\prime}=f(H^{\prime})$, and hence $\deg f|_{H^{\prime}}=m=\deg f$. Namely, $H^{\prime}\in S(f)$.
(4) The “only if” part follows from (3). Conversely, suppose $t_{y}=1$ and $L_{Y}\geq 0$. Thus $H_{X}^{\prime}:=f^{*}L_{Y}$ is in $|H_{X}|$ and hence it and then also $L_{Y}$ are reduced and irreducible, so $L_{Y}=f(H_{X}^{\prime})$. By the projection formula, $mL_{Y}=f_{*}H_{X}^{\prime}$, so $H_{X}^{\prime}\in S(f)$.
(5) Assume first ${\rm Codim}\,S(f)\leq 1$ in $|H_{X}|$. By (3), we have $f^{*}L_{Y}\sim H_{X}$. Taking top intersection and by assumption, $d=H_{X}^{n}=(f^{*}L_{Y})^{n}=mL_{Y}^{n}\geq dL_{Y}^{n}$. Thus $L_{Y}^{n}\leq 1$. Since $S(f)=f^{*}|L_{Y}|$ (cf. (3)), we have $\dim|L_{Y}|=\dim S(f)\geq\dim|H_{X}|-1=n$. Thus
(*) $\dim H^{0}(Y,L_{Y})\geq n+1\geq n+L_{Y}^{n}.$
Claim 4.3.
$|L_{Y}|$, or equivalently $S(f)$, defines a generically finite map $\Phi_{|L_{Y}|}$ (resp. $\Phi_{S}$).
If $S(f)=|H_{X}|$, then as $H_{X}$ is very ample, the claim is true. If $\dim S(f)=\dim|H_{X}|-1$, the linearity of $S(f)=f^{*}|L_{Y}|$ and Proposition 2.2 (2) imply that $S(f)$ is the restriction to $X$ of a sub-linear system of $|\mathcal{O}_{{\mathbb{P}}^{n+1}}(1)|$ defined by one linear equation in the dual space $({\mathbb{P}}^{n+1})^{*}$. Thus $S(f)=\{H_{P}\in|{\mathcal{O}}_{{\mathbb{P}}^{n+1}}(1)|~{};~{}Q\in H_{P}\}|_{X}$ for some point $Q\in{\mathbb{P}}^{n+1}$. Hence the map $\Phi_{S}$ is the projection $\varphi:{\mathbb{P}}^{n+1}\setminus\{Q\}\to{\mathbb{P}}^{n}$ restricted to $X$. If $\varphi(X)={\mathbb{P}}^{n}$ then the claim follows. Otherwise, $\varphi(X)$ is a hypersurface in ${\mathbb{P}}^{n}$ in coordinates $X_{0},\dots,X_{n}$ (while ${\mathbb{P}}^{n+1}\supset X$ is in coordinates $X_{0},\dots,X_{n+1}$) so that $X$ is the cone over $\varphi(X)$ and hence defined as $X=V(F(X_{0},\dots,X_{n}))$ for some homogeneous polynomial $F$ in the first $n+1$ coordinates of ${\mathbb{P}}^{n+1}$. This contradicts $X$ being a general hypersurface. The claim is proved.
We return back to the proof of the proposition. With Claim 4.3 and the ()(*)( ∗ ) above, Proposition 4.1 implies $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$. Also $d=H_{X}^{n}=(f^{*}L_{Y})^{n}=\deg f$.
Conversely, if $m=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$, then $t_{y}^{n}d=(t_{y}H_{X})^{n}=(f^{*}L_{Y})^{n}=m$, so $t_{y}=1$. Thus, by (4) (3), $S(f)=f^{*}|L_{Y}|$ and $\dim S(f)=\dim H^{0}(Y,L_{Y})-1=n=\dim|H_{X}|-1$.
This proves the proposition. ∎
Remark 4.4.
In the case of Proposition 4.2 (5), if we assume $d\geq 2n+3$ then Lemma 5.4 (whose assumption is satisfied because $\deg f=d$ as in the proof of Propostion 4.2) shows that the fiber $Z_{y}$ for a general point $y\in Y$ of $f$ lies on the unique line $\ell_{y}$ in ${\mathbb{P}}^{n+1}$. In fact [5, Lemma 2.4] says that $f^{-1}(y)$ lies on the unique line $\ell_{y}$ if all points in $f^{-1}(y)$ are distinct points. Then imitating [5, Proof of Lemma 4.1] and [6, Proof of Theorem C], one might be able to show that the morphism $f:X\to Y$ is nothing but a projection from a point in ${\mathbb{P}}^{n+1}\setminus X$. However, we do not pursue a proof here, since this paper focuses on the characterization of $Y$.
When $\deg f=p$ is prime, there is an upper bound for $\deg f$ which depends on $\deg X=d$ linearly.
Theorem 4.5.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d$ and $f:X\to Y$ a surjective morphism of prime degree to a normal projective variety $Y$ with at worst $\varepsilon$-lc singularities. Assume either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$. Assume further either $Y$ is factorial (or just $L_{Y}^{n}$ is an integer) or $-K_{Y}$ is ample, when $f^{*}L_{Y}\sim H_{X}$. Then:
  1. (1)
    There are constants $D_{j}=D_{j}(n,\varepsilon)$ with $D_{1}\geq 1$, with $D_{1}=1$ when $Y$ is factorial (or just $L_{Y}^{n}$ is an integer) and with $D_{2}=(D_{2}^{\prime})^{1/(n-1)}$ in terms of upper bound $D_{2}^{\prime}=D_{2}^{\prime}(n,\varepsilon)$ of $(-K_{Y})^{n}$ (as in Birkar [8, Theorem 1.1], or [27, Corollary 1] with $D_{2}^{\prime}=\{n(n+1)\}^{n}$ when $Y$ is smooth), such that
    $\deg f\,\leq\,{\rm max}\,\{D_{1}d,\,\,\,\,\frac{D_{2}}{d^{1/(n-1)}}\}.$
  2. (2)
    If $Y$ is factorial (or just $L_{Y}^{n}$ is an integer) and $d\geq D_{2}^{(n-1)/n}$, then $\deg f\leq d$.
Proof.
Note that (2) follows from (1). We prove (1). Let $p=\deg f$ which is assumed to be a prime number. We may assume $p\geq d$. By our assumption, all conditions in Proposition 2.2 are satisfied. So $Y$ is ${\mathbb{Q}}$-factorial and klt with $\rho(Y)=1$, and $\operatorname{Pic}(X)={\mathbb{Z}}[H_{X}]$ with $H_{X}$ a general section of ${\mathcal{O}}_{{\mathbb{P}}^{n+1}}(1)|_{X}$; every member of $|H_{X}|$ is reduced and irreducible.
Consider $S(f):=\{H\in|H_{X}|\,;\,\deg f|_{H}=\deg f\}$. If $\dim S(f)\geq n$, then $p=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$, by Proposition 4.2. Thus we may assume $\dim S(f)\leq n-1$, i.e., $S(f)$ is of codimension $\geq 2$ in $|H_{X}|$.
Let $H_{Y}:=f(H_{X})$, an integral and ${\mathbb{Q}}$-Cartier ample divisor. Let $f_{H_{X}}=f|_{H_{X}}:H_{X}\to H_{Y}$ and $m_{H_{X}}=\deg f_{H_{X}}$. Write $f^{*}H_{Y}\sim b_{H_{X}}H_{X}$. By the projection formula, $pH_{Y}=f_{*}f^{*}H_{Y}\sim b_{H_{X}}m_{H_{X}}H_{Y}$. So $p=b_{H_{X}}m_{H_{X}}$. Since $p$ is prime, $(m_{H_{X}},b_{H_{X}})=(p,1)$ or $(1,p)$. Since we assume $S(f)$ has codimension $\geq 2$ in $|H_{X}|$, we have $(m_{H_{X}},b_{H_{X}})=(1,p)$ for general $H_{X}$. Namely, $f_{*}H_{X}=H_{Y}$ and $f^{*}H_{Y}\sim pH_{X}$. Write $H_{Y}\sim h_{y}L_{Y}$. Then $h_{y}t_{y}H_{X}\sim f^{*}(h_{y}L_{Y})=f^{*}H_{Y}\sim pH_{X}$. Hence $h_{y}t_{y}=p$. So $(h_{y},t_{y})=(p,1)$ or $(1,p)$.
Consider the case $(h_{y},t_{y})=(p,1)$. Then $H_{Y}\sim pL_{Y}$ and $f^{*}L_{Y}\sim H_{X}$. If $L_{Y}^{n}$ is an integer (so $\geq 1$, since $L_{Y}$ is ample), we have
$d=H_{X}^{n}=(f^{*}L_{Y})^{n}=pL_{Y}^{n}\geq p.$
If $-K_{Y}$ is ample, by Proposition 4.2 (2), we have
$p\leq D_{1}d$
for some constant $D_{1}=D_{1}(n,\varepsilon)$.
Consider the case $(h_{y},t_{y})=(1,p)$. Then $H_{Y}\sim L_{Y}$ and $f^{*}L_{Y}\sim pH_{X}$.
Claim 4.6.
$-K_{Y}$ is ample.
We prove the claim. Suppose the contrary that $-K_{Y}$ is not ample. Then by Theorem 1.2, $K_{Y}$ is ample. By the ramification divisor formula, $(d-n-2)H_{X}\sim K_{X}=f^{*}K_{Y}+R_{f}$. Pushing forward using $f_{*}$, we get $(d-n-2)H_{Y}\geq pK_{Y}\geq pL_{Y}\sim pH_{Y}$. So $p\leq d-n-2$, contradicting the assumption $p\geq d$. This proves the claim.
We return back to the proof of the theorem. Write $-K_{Y}\sim rL_{Y}\sim rH_{Y}$ for some integer $r\geq 1$. Then the ampleness of them, imply
(*) $p^{n}d=(pH_{X})^{n}=(f^{*}H_{Y})^{n}=pH_{Y}^{n}\leq p(-K_{Y})^{n}\leq pD_{2}^{\prime}(n,\varepsilon)$
for some constant $D_{2}^{\prime}=D_{2}^{\prime}(n,\varepsilon)$ as specified in the statement of the theorem. Now (1) follows from the ()(*)( ∗ ) above by letting $D_{2}=(D_{2}^{\prime})^{1/(n-1)}$. ∎
Proof of Theorem 1.3 and Remark 1.5 (4).
Theorem 1.3 (1) follows Proposition 4.2 (4) and (2). For Theorem 1.3 (2), by Theorem 1.2, we may assume that $\deg f\geq C(n,d)$ and hence $-K_{Y}$ is ample. Then it follows by taking the current $D_{2}$ to be the maximum of its name sake in Theorem 4.5 and $C(n,d)$. For Theorem 1.3 (3) and Remark 1.5 (4), we may assume $\deg f\geq d$, and then they follow from Proposition 4.2 (5). ∎
Proof of Theorem 1.4 and Remark 1.5 (5).
We assume either $n\geq 3$ and $d\geq 3$, or $n=2$ and $d\geq 4$ (instead of $n\geq 3$ and $d\geq 3$). Theorem 1.4 (1) follows Proposition 4.2 (4) and (2). Theorem 1.4 (2) implies (3), while (2) itself follows from Theorem 4.5 with $\varepsilon=1$ since $Y$ is klt and factorial and hence has at worst canonical singularities. ∎
As consequences of Theorem 4.5, we now give sufficient conditions for the target $Y$ of a very general hypersurface $X\subset{\mathbb{P}}^{n+1}$ to be ${\mathbb{P}}^{n}$. We start with dimension $3$ and then dimension $\geq 4$.
Let $Y$ be a smooth Fano $n$-fold with $\operatorname{Pic}Y={\mathbb{Z}}[H]$. Write $-K_{Y}\sim rH$. We call $r$ the Fano index of $Y$ and $n+1-r$ the Fano co-index of $Y$. It is known that $r\leq n+1$, and $r=n+1$ (resp. $r=n$) holds if and only if $Y\cong{\mathbb{P}}^{n}$ (resp. $Y$ is a hyperquadric in ${\mathbb{P}}^{n+1})$; see [24, Corollary to Theorem 1.1, Corollary to Theorem 2.1].
Corollary 4.7.
In Theorem 1.4, assume further
  1. (1)
    $\deg f$ is prime,
  2. (2)
    $Y$ is smooth,
  3. (3)
    $\deg f\geq d$,
  4. (4)
    $n=3$, and
  5. (5)
    $d>n(n+1)=12$.
Then $\deg f=d$ and $Y\cong{\mathbb{P}}^{3}$.
Proof.
Since $Y$ is smooth, $\operatorname{Pic}Y={\mathbb{Z}}[L_{Y}]$ by Propositions 4.2 and 2.2. Let $p=\deg f$, a prime number. If $\dim S(f)\geq n$, then $p=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$, by Proposition 4.2. Thus, we may assume $\dim S(f)\leq n-1=2$. By Theorem 1.4 and Remark 1.5 (2), we have $\deg f\leq d$, so $\deg f=p=d$. Since $\dim S(f)\leq n-1$, by the proof of Theorem 4.5, we may assume that $(m_{H_{X}},b_{H_{X}})=(1,p)$ for general $H_{X}$, i.e., $f_{*}H_{X}=H_{Y}:=f(H_{X})$ and $f^{*}H_{Y}\sim pH_{X}$; further, $(h_{y},t_{y})=(p,1)$ or $(1,p)$, i.e., $H_{Y}\sim pL_{Y}$ and $f^{*}L_{Y}\sim H_{X}$, or $H_{Y}\sim L_{Y}$ and $f^{*}L_{Y}\sim pH_{X}$.
By Theorem 1.2, $K_{Y}$ or $-K_{Y}$ is ample. Thus $-K_{Y}$ is ample (cf. [3, Proof of Lemma 3.0]).
Suppose $H_{Y}\sim L_{Y}$ and $f^{*}L_{Y}\sim pH_{X}$. Then, $p^{n+1}=p^{n}d=(f^{*}L_{Y})^{n}=pL_{Y}^{n}$. Also, by [27, Corollary 1], we have (with $n=3$)
$(n(n+1))^{n}\geq(-K_{Y})^{n}\geq L_{Y}^{n}=p^{n}=d^{n}$
contradicting our assumption.
Thus, $H_{Y}\sim pL_{Y}$ and $f^{*}L_{Y}\sim H_{X}$, hence $L_{Y}^{3}=1$. Write $-K_{Y}\sim rL_{Y}$. We may assume $r\leq n=3$. Then $(-K_{Y})^{3}=r^{3}$. Using the classification theory of Fano threefolds with Picard number $\rho=1$ (cf. [35, Table 12.2]), we have $r=2$ and $\dim H^{1,2}(Y)=21$ (so $Y$ is a del Pezzo threefold of degree 1). Note that the Hodge pair $(h^{3,0},h^{1,2})$ of $Y$ is $(0,21)$ while that of $X$ is (+,)(+,*)( + , ∗ ) (because $d>n+1$) and of ${\mathbb{P}}^{3}$ is $(0,0)$. We get a contradiction to [3, Corollary 2.2]. ∎
Corollary 4.8.
In Theorem 1.4, assume further
  1. (1)
    $\deg f$ is prime,
  2. (2)
    $Y$ is a smooth Fano variety,
  3. (3)
    $\deg f\geq d$,
  4. (4)
    $n\geq 4$,
  5. (5)
    $d>n(n+1)$, and
  6. (6)
    Fano index of $Y$ is $\geq n-2$ and $\neq n-1$.
Then $\deg f=d$ and $Y\cong{\mathbb{P}}^{n}$.
Proof.
Since $Y$ is smooth, $\operatorname{Pic}Y={\mathbb{Z}}[L_{Y}]$ by Propositions 4.2 and 2.2. Let $p=\deg f$, a prime number. If $\dim S(f)\geq n$, then $p=d$, $Y={\mathbb{P}}^{n}$ and ${\mathcal{O}}(L_{Y})={\mathcal{O}}(1)$, by Proposition 4.2. Thus we may assume $\dim S(f)\leq n-1$. By the same argument as in Corollary 4.7, we may assume $p=d$, $H_{Y}\sim pL_{Y}$, $f^{*}L_{Y}\sim H_{X}$ (so $t_{y}=1$), and $L_{Y}^{n}=1$. On the other hand, by Ambro [1, Main Theorem], we have $\dim|L_{Y}|\geq n-1$. Now $S(f)=f^{*}|L_{Y}|$ by Proposition 4.2. Therefore, $\dim S(f)=\dim|L_{Y}|=n-1$. So Fujita’s $\Delta$-genus of $(Y,L_{Y})$ is $1$, which would reach a contradiction.
Indeed, by applying Fujita [13, Theorem 2.1 and Proposition 3.2], a general member $D$ in $|L_{Y}|$ is irreducible and smooth. The polarized pair $(Y,L_{Y})$ satisfies the following:
  1. (1)
    $Y$ is a smooth Fano variety with Fano index $\geq n-2$ and $\neq n-1$, i.e., of Fano co-index $\leq 3$ but $2$, and with $\operatorname{Pic}Y={\mathbb{Z}}[L_{Y}]$,
  2. (2)
    $L_{Y}^{n}=1$,
  3. (3)
    $\dim|L_{Y}|=n-1$,
  4. (4)
    Fujita’s $\Delta$-genus $\Delta(Y,L_{Y})=1$, and
  5. (5)
    $\operatorname{Pic}D=1$ for a general member $D$ in $|L_{Y}|$ (cf. [28, Example 3.1.25]).
Then the pair $(D,L_{Y}|_{D})$ satisfies again all the above conditions (1)–(5). Indeed, by the adjunction formula, $D$ is a smooth Fano variety of dimension $n-1$ and Fano index $\geq(n-1)-2$. Since $Y$ is a smooth Fano variety, $H^{1}(Y,{\mathcal{O}}_{Y})=0$ by Kodaira vanishing. Therefore, we keep
$\Delta(D,L_{Y}|_{D})=\Delta(Y,L_{Y})=1$
by [13, (1.5)]. And by [28, Example 3.1.25], the restriction $\operatorname{Pic}(Y)\to\operatorname{Pic}(D)$ is an isomorphism because $n\geq 4$. Therefore $L_{Y}|_{D}$ is an ample generator of $\operatorname{Pic}D$ and $(L_{Y}|_{D})^{n-1}=1$. Thus $Y$ and $D$ share the same Fano co-index. Also from the exact sequence
$0\to{\mathcal{O}}_{Y}\to{\mathcal{O}}_{Y}(L_{Y})\to{\mathcal{O}}_{D}(L_{Y}|_{D})\to 0,$
we have $\dim|L_{Y}|_{D}|=\dim|L_{Y}|-1$ because $H^{1}(Y,{\mathcal{O}}_{Y})=0$. It gives $\dim|L_{Y}|_{D}|=n-2$. Then we can apply again [13, Theorem 2.1 and Proposition 3.2].
This process reduces to the case where $\dim Y=4$. We note that this process keeps the Fano co-index. Then we have a smooth Fano threefold $D\in|L_{Y}|$ again by [13, Theorem 2.1 and Proposition 3.2]. This $D$ satisfies the above conditions (1)–(4). As in Corollary 4.7, write $-K_{D}\sim rL_{Y}|_{D}$ with $r\leq\dim D+1=4$, and use $(-K_{D})^{3}=r^{3}$ and [35, Table 12.2]. We see that either $r=4$ and $D={\mathbb{P}}^{3}$, or $r=2$. If $D={\mathbb{P}}^{3}$, then ($1=$) $\Delta(D,L_{Y}|_{D})=\Delta({\mathbb{P}}^{3},{\mathcal{O}}(1))=0$, a contradiction. If $r=2$, then $D$ and hence $Y$ have Fano co-index 2, which we exclude by the assumption. This proves the corollary. ∎
5. More conjectures and relations among them
In this section we will discuss a related conjecture about dominant rational maps. Precisely, one may also consider the following one about dominant rational maps from a very general hypersurface (cf. Guerra and Pirola in [17, Conjecture, page 290]):
Conjecture 5.1.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d\geq n+3$, and $Y$ a smooth projective $n$-fold. Suppose there is a dominant rational map $f:X\dashrightarrow Y$ with ${\rm deg}\,f>1$. Then $Y$ is rationally connected.
By Hodge theory (cf. Guerra and Pirola [17, Proposition 3.5.2]), to prove Conjecture 5.1, it suffices to consider dominant rational maps $f:X\dashrightarrow\widetilde{Y}$, where $\widetilde{Y}$ is a smooth model of $Y$, which is simply connected and without holomorphic global $n$-forms, i.e., $p_{g}(\widetilde{Y})=0$.
In [17], Guerra and Pirola proved that the $Y$ in Conjecture 5.1 is not of general type for the case $n=2$ and $5\leq d\leq 11$. Conjecture 5.1 for $n=2$ is proved in the joint paper [29] of Pirola and the first-named author of the present paper. In Conjecture 5.1, the assumption ‘general’ is clearly necessary. Indeed, if one chooses a special surface $X$ in ${\mathbb{P}}^{3}$, it might admit a dominant rational map to a surface of general type $Y$. Classical Godeaux surfaces, which are of general type, are obtained as ${\mathbb{Z}}/(5)$-quotients of ${\mathbb{Z}}/(5)$-invariant quintics (cf. Godeaux [16]). Also for instance, a Fermat hypersurface of degree $d$ in $\mathbb{P}^{n+1}$ with $d=d_{1}a$ can map to a Fermat hypersurface of degree $d_{1}$ in $\mathbb{P}^{n+1}$, which is not rationally connected when $d_{1}\geq n+2$, by taking $a$-th power.
In [29], the assumption ‘very general’ is used to deal with the case $n=2$ and to exclude countably many different deformation types of non-ruled simply connected minimal surfaces with $p_{g}=0$.
Here, a projective variety is rationally connected in the sense of Campana and Kollár-Miyaoka-Mori, if two general points are connected by a rational curve; this is a birational notion (cf. Kollár [25, VI, Proposition 3.3]). We consider ‘rational connectedness’ because this assumption fits well with the birationality and induction argument on the dimension.
We remark that one approach to prove Conjecture 5.1 is by induction via the following conjecture and Proposition 5.3. Indeed, Conjecture 5.1 in dimension $\leq n-1$ and Conjecture 5.2 in dimension $n$ imply Conjecture 5.1 in dimension $n$.
Conjecture 5.2.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d\geq n+3$, and $W$ a smooth projective $n$-fold which is non-uniruled. Let $H_{X}:=X\cap\,H$, where $H$ is a very general hyperplane in ${\mathbb{P}}^{n+1}$. Suppose there is a dominant rational map $h:X\dashrightarrow W$ with ${\rm deg}\,h>1$. Then the restriction of $h$ to $H_{X}$ is a non-birational map from $H_{X}$ to its image in $W$.
We remark that although there are countably many different deformation types for simply connected minimal non-ruled surfaces with $p_{g}=0$, the dimension of each moduli space of simply connected minimal non-ruled surfaces with $p_{g}=0$ is bounded by 19 (cf. [17, Corollary 2.5.3] for surfaces of general type, [29, Proof of Theorem 1.1] for surfaces with Kodaira dimension $\kappa=1$). This helps to solve the case $n=2$ for Conjectures 5.1 and 1.1 in [29]. However, when $\dim Y\geq 3$, there are no known results for the dimension of moduli of simply connected minimal non-ruled varieties with $h^{i}(Y,{\mathcal{O}}_{Y})=0$ for all $i>0$.
Proposition 5.3.
Let $g:X\dasharrow Z$ be a dominant rational map between projective varieties over an algebraically closed field of characteristic zero, where $X$ is normal and $\dim\,X>\dim\,Z\geq 1$. Let $H$ be a nef and big Cartier divisor, and $\Lambda$ a non-trivial linear subsystem of $|H|$ with ${\rm Bs}\,\Lambda=\emptyset$. Then for general $H_{1},\dots,H_{r}$ in $\Lambda$ with $r=\dim\,X-\dim\,Z$, the map $g$ restricts to the intersection $H_{1}\cap\ldots\cap H_{r}$ as a generically finite dominant map.
Proof.
Replacing $Z$ and $X$ by smooth models $Z^{\prime}$ and $X^{\prime}$, and pulling back $H$ and $\Lambda$ to $X^{\prime}$, we may assume that $X$ and $Z$ are smooth and $g$ is a surjective morphism.
Since the Iitaka dimension $\kappa(X,H)=\dim\,X>\dim\,Z$, any $H^{\prime}$ in $\Lambda$ dominates $Z$, or else $H^{\prime}\leq f^{*}D$ for some divisor $D$ on $Z$ and
$\kappa(X,H)=\kappa(X,H^{\prime})\leq\kappa(X,f^{*}D)=\kappa(Z,D)\leq\dim\,Z,$
which is a contradiction. By Bertini’s theorem, a general member $H_{1}$ of $\Lambda$ is irreducible and smooth as $\Lambda$ is base point free and $\dim X\geq 2$. Replace $X$ by $H_{1}$ and $\Lambda$ by $\Lambda\,|_{H_{1}}$, which remains a non-trivial base point free linear subsystem of $|H_{1}|_{H}|$ with $H_{1}|_{H}$ being nef and big. By induction, we have a surjective morphism
$H_{1}\cap\ldots\cap H_{r}\to Z$
between varieties of the same dimension, thereby proving the proposition. ∎
To end this section, we remark that there is a lower bound for $\mathrm{deg}~{}f$ if the generically finite map $f:X\dashrightarrow Y$ is a non-birational dominant map. Lemma 5.4 below is essentially from [5].
Lemma 5.4.
Let $X\subset{\mathbb{P}}^{n+1}$ be a very general hypersurface of degree $d\geq n+3$ with $n\geq 2$, and $Y$ a smooth projective $n$-fold. Suppose there is a dominant rational map $f:X\dashrightarrow Y$ with $\mathrm{deg}~{}f>1$. Then the following assertions hold.
  • (1)
    $\mathrm{deg}~{}f\geq d-n$.
  • (2)
    If $\mathrm{deg}~{}f\leq 2d-2n-3$, then the fiber $Z_{y}$ for a general point $y\in Y$ of $f$ lies on a line $\ell_{y}$ in ${\mathbb{P}}^{n+1}$ and $\mathrm{deg}~{}f\leq d$.
Proof.
By Claim 2.4, $p_{g}(Y)=0$. Therefore, given $\omega\in H^{0}(X,{\mathcal{O}}(K_{X}))$, its trace (cf. [5, Definition 2.1]) in $H^{0}(Y,{\mathcal{O}}(K_{Y}))$ vanishes. Accordingly, the general fiber of $f$ has the Cayley-Bacharach property (cf. Bastianelli-Cortini-De Poi [5, Definition 2.2]) with respect to ${\mathcal{O}}(K_{X})={\mathcal{O}}_{X}(d-n-2)$. Namely, if a section of ${\mathcal{O}}(K_{X})$ vanishes at all but one point of the general fiber then it vanishes at every point of the general fiber. Hence we have $\mathrm{deg}~{}f\geq d-n$; see [5, Proposition 2.3, Lemma 2.4] or [6, Theorem 2.3]. This proves (1).
For (2), suppose that $\mathrm{deg}~{}f\leq 2(d-n-2)+1$. The rational map $f$ gives a correspondence $Z\subseteq X\times Y$ with vanishing trace map from $H^{0}(X,{\mathcal{O}}(K_{X}))$ to $H^{0}(Y,{\mathcal{O}}(K_{Y}))$. By [5, Lemma 2.4 or Theorem 2.5], general fiber $Z_{y}$ of $f$ lies on a line $\ell_{y}\subset{\mathbb{P}}^{n+1}$. Since $\deg X=d$, we have $\mathrm{deg}~{}f\leq d$ because $Z_{y}$ lies on a line $\ell_{y}$. ∎